OpenStudy (anonymous):

Simplify:

OpenStudy (anonymous):

\[7\sqrt{3} + 8\sqrt{3} - 2\sqrt{2}\]

OpenStudy (anonymous):

\[\huge \sqrt{3}(7+8)-2\sqrt{2} \rightarrow \sqrt{3}(15)-2\sqrt{2} \]